# Oldies but Goldies: The Potential of Character N-grams for Romanian Texts

### v1

@@ Abstract
This study addresses the problem of authorship attribution for Romanian texts using the ROST corpus, a standard benchmark in the field. We systematically evaluate six machine learning techniques—Support Vector Machine (SVM), Logistic Regression (LR), k-Nearest Neighbors (k-NN), Decision Trees (DT), Random Forests (RF), and Artificial Neural Networks (ANN)—employing character n-gram features for classification. Among these, the ANN model achieved the highest performance, including perfect classification in four out of fifteen runs when using 5-gram features. These results demonstrate that lightweight, interpretable character n-gram approaches can deliver state-of-the-art accuracy for Romanian authorship attribution, rivaling more complex methods. Our findings highlight the potential of simple stylometric features in resource-constrained or under-studied language settings.
@@ 1 Introduction
  • Preprocessing steps: such as text normalization and punctuation handling [2];
@@
  • Computational methods: ranging from traditional statistical techniques (e.g., stylometry, Principal Component Analysis) to machine learning algorithms (e.g., Support Vector Machines, Random Forests), deep learning models (e.g., Recurrent Neural Networks, transformers), and modern large language model (LLM) approaches [6];
@@
  • class imbalance (disproportionate representation of authors or genres),
  • generalizability issues (e.g., models failing on cross-domain or cross-genre texts).
@@
Dataset partitioning into training, testing, and validation subsets also affects performance outcomes, as different splits can yield varying results due to random sampling effects. To mitigate this, researchers commonly employ multiple splits or cross-validation techniques to obtain more reliable performance estimates.
@@ 2.2 N-gram Features in Authorship Attribution
N-grams are contiguous sequences of $N$ items, such as characters, words, or other tokens, extracted from text to capture patterns that may be indicative of an author’s unique style. Typically, n-grams are constructed at the character, word, or syntactic level [5].
@@ 2.2.1 Advantages of Character N-grams
  • Robustness: Character n-grams are resilient to infrequent errors such as grammatical mistakes or punctuation slips because their discriminative power derives primarily from frequent, recurring patterns [12].
@@
  • Computational Efficiency: Extracting character n-grams is straightforward and computationally inexpensive, as it involves direct processing of raw text without the need for complex preprocessing.
@@ 2.3 Computational Methods
A variety of classification algorithms [16, 17, 18, 19, 20, 21] have been applied to authorship attribution (AA) using character n-gram features. Below, we summarize the key methods employed in this study:
@@ 2.4.1 Accuracy
Accuracy measures the proportion of correctly classified texts across all classes:
Accuracy=Total Correct PredictionsTotal Number of $\text{Accuracy}=\frac{\text{Total Correct Predictions}}{\text{Total Number of Samples}}$
While widely used, accuracy can be misleading for imbalanced datasets, as it may overemphasize performance on majority classes.
@@ 3.1 The dataset
  • Avram (2023) leveraged a Romanian-specific BERT model (dumitrescustefan/bert-base-romanian-cased-v1), achieving approximately 87% macro-accuracy [9].
@@ 3.2 Data Preprocessing
  • Punctuation Preservation: All other special characters (punctuation marks) were left unchanged, as patterns in punctuation usage may reflect individual authorial style and thus serve as valuable features.
@@ 3.3.2 Vectorization via TF-IDF
  • Cell values = TF-IDF weights for corresponding N-gram/document pairs
@@ 3.4.1 Implementation Details
  • Feature Extraction: TfidfVectorizer from scikit-learn for generating character N-gram feature representations.
@@ 3.4.2 Classification Algorithms
We evaluated a diverse set of supervised learning algorithms, all implemented via the scikit-learn library:
  1. 1.
    Support vector machine (SVM)
  2. 2.
    Logistic Regression (LR)
  3. 3.
    k-Nearest Neighbor (k-NN)
  4. 4.
    Decision trees (DT)
  5. 5.
    Random forest (RF)
  6. 6.
    Artificial neural networks (ANN)
@@ 3.4.3 Model Evaluation
For model assessment, we use standard metrics including accuracy, macro-accuracy (aka. balanced accuracy or macro-averaged accuracy), and the classification report (precision, recall, F1-score), as provided by scikit-learn.
@@ 3.5 Experimental Procedure
For each classification method described previously, we conducted a series of experiments to identify the configuration yielding the best performance. Our investigation included the following considerations:
  • Parameter Exploration: For the k-Nearest Neighbors (k-NN) algorithm, we evaluated five different values of k to determine the optimal neighborhood size.
  • Robustness to Randomness: To mitigate the effects of stochastic variability inherent in certain models, we repeated the training and evaluation of Decision Trees (DT), Random Forests (RF), and Artificial Neural Networks (ANN) five times, each with a different random seed. This procedure allowed us to assess the stability and reliability of the results.
@@ 3.6 Parameter Tuning and Model Selection
k-NN metric: minkowski n_neighbors = $\{3,5,7,9,11\}$
DT criterion: gini random_state = $\{7,17,42,67,101\}$
RF n_estimators: 100, criterion: gini random_state = $\{7,17,42,67,101\}$
ANN activation: relu random_state = $\{7,17,42,67,101\}$
@@
Table 1: Summary of machine learning models and their hyperparameters. To ensure that performance differences reflect model characteristics rather than random variation, Decision Trees (DT), Random Forests (RF), and Artificial Neural Networks (ANN) were each trained and evaluated five times using different random seeds. For k-Nearest Neighbors (k-NN), five different values of the number of neighbors were tested.
@@ 3.7.1 Impact of Letter Casing on Classification Performance
Figure 1 depicts the variation in MAcc differences across all classifiers and n-gram sizes. Overall, lowercase texts tend to yield marginally better results, particularly for $k>3$ and for the k-NN and Logistic Regression (LR) classifiers. However, in most cases, the difference remains below 0.02, which is negligible and likely falls within the range of normal variation due to hyperparameter tuning.
For the Artificial Neural Network (ANN) model, the average difference between casing strategies is even smaller, approximately 0.009. Interestingly, at n-gram size $N=5$, the trend reverses slightly, with original-case texts performing better, as further illustrated in Figure 4.
@@ 3.7.2 The Impact of N-gram Size on Classification Performance
A noticeable decline in MAcc is observed for the Random Forest (RF) classifier (Figure 2e) as the N-gram size increases. Conversely, the Artificial Neural Network (ANN) shows a general upward trend in MAcc with larger N-gram dimensions. For Support Vector Machines (SVM), the highest MAcc is achieved at $k=4$, while k-Nearest Neighbors (k-NN) attains its peak performance at $k=2$. The remaining methods do not exhibit a clear monotonic trend in MAcc relative to the N-gram size.
Figure 2: Variation of Macro-Accuracy (MAcc) across different N-gram sizes ($N=2$ to $5$) for all six classification methods: (a) SVM, (b) LR, (c) k-NN, (d) DT, (e) RF, and (f) ANN. This analysis examines how varying the N-gram size affects the models’ performance. Values shown represent the average of the macro-accuracy scores computed over multiple experimental runs.
@@ 3.7.3 Best Results
The Artificial Neural Network (ANN) model achieved the best overall performance, with average accuracy and macro-accuracy (MAcc) values exceeding 0.9, as summarized in Table 2. Tables 3 and 4 present the average accuracy and MAcc for each model across N-gram sizes ($N=2$ to $5$), reported separately for original-case and lowercase texts.
@@
ANN consistently outperformed other models, except at $k=2$, where the Random Forest (RF) classifier achieved superior results. This trend is illustrated graphically in Figure 3, which visualizes the MAcc values from the aforementioned tables.
Further details on ANN performance are provided in Tables 7 and 8, where for each $k$ and data split (with five runs per split), we report the average, standard deviation, and best values for both accuracy and macro-accuracy.
For ANN, the highest performance was observed at $k=4$ using original-case letters, and at $k=5$ using lowercase letters—with lowercase slightly outperforming original case by 0.003 in MAcc. Conversely, RF achieved the best results at $k=2$, with average macro-accuracy values of 0.914 (original case) and 0.919 (lowercase).
Examining RF results further (Figures 2e and Tables 56 from appendix A), we observe a slight decrease in average accuracy as $k$ increases, although MAcc values across different $k$ are not clearly separated.
Notably, ANN achieved perfect classification (accuracy = 1.0) for one test, as shown in Tables 78 from appendix A). To further investigate this, we extended the experiments to 15 runs with different random seeds:
@@
for both original-case and lowercase texts. As shown in Figure 4, perfect accuracy was reached in four cases: twice for $random\_state=42$ (in both casing conditions) and two additional times for original-case texts. The standard deviation of MAcc across these runs was approximately 0.02 (0.02039 for lowercase, 0.01938 for original case, and 0.02040 overall). No clear pattern emerged relating accuracy variation to the random seed. However, original-case texts tended to perform slightly better on average, with a 0.01 higher MAcc (0.974 vs. 0.962 for lowercase), reversing the trend observed in Tables 3 and 4. Given the small magnitude of these differences, and the reversed trend in the second set of experiments for ANN, we conclude that casing (original vs. lowercase) does not have a significant impact on classification performance.
@@ 3.8 Comparison with Existing Results
Our approach, employing an Artificial Neural Network (ANN), achieved an initial average macro-accuracy of 0.935, with a perfect classification (macro-accuracy = 1.0) observed in one of five data splits. Extending the evaluation to 15 runs with varying random seeds (see Section 3.7.3) improved the average macro-accuracy to 0.974 for original-case texts, with a standard deviation of approximately 0.02. Perfect classification was attained in four runs: once using lowercase texts and three times using original case. Interestingly, for the random seed 42, perfect classification was achieved in both casing conditions.
These results demonstrate that our lightweight, character N-gram based approach with ANN matches or surpasses existing benchmarks for Romanian authorship attribution, offering a competitive alternative to more complex models such as BERT and/or transformer-based approaches.
@@ 4 Conclusions and Future Work
Appendix A Detailed performance results achieved using Random Forest (RF) and Artificial Neural Networks (ANN) methods
N-gram Split Macro-Accuracy Accuracy
N no. avg std best avg std best
2 1 0.886 0.027 0.913 0.879 0.021 0.901
2 2 0.915 0.008 0.922 0.911 0.009 0.926
2 3 0.920 0.038 0.963 0.921 0.028 0.951
2 4 0.933 0.013 0.955 0.928 0.014 0.938
2 5 0.915 0.025 0.954 0.921 0.020 0.951
3 1 0.851 0.031 0.890 0.842 0.029 0.877
3 2 0.916 0.035 0.955 0.923 0.031 0.963
3 3 0.889 0.020 0.911 0.909 0.020 0.938
3 4 0.919 0.016 0.939 0.921 0.013 0.938
3 5 0.875 0.017 0.892 0.894 0.010 0.901
4 1 0.778 0.027 0.812 0.770 0.024 0.802
4 2 0.865 0.009 0.878 0.884 0.010 0.901
4 3 0.839 0.028 0.874 0.877 0.028 0.914
4 4 0.895 0.029 0.949 0.899 0.021 0.938
4 5 0.865 0.017 0.880 0.884 0.018 0.901
5 1 0.763 0.017 0.788 0.756 0.016 0.778
5 2 0.864 0.018 0.885 0.889 0.016 0.914
5 3 0.834 0.006 0.844 0.886 0.005 0.889
5 4 0.867 0.044 0.937 0.877 0.032 0.926
5 5 0.879 0.011 0.893 0.906 0.013 0.926
Table 5: The average (avg), standard deviation (std), and best scores for Macro-Accuracy and Accuracy achieved by the Random Forest (RF) classifier over five runs, evaluated for each N-gram size ($N=2$ to $5$) and across multiple data splits, and for the original letter casing of the text.
N-gram Split Macro-Accuracy Accuracy
N no. avg std best avg std best
2 1 0.887 0.020 0.910 0.886 0.020 0.914
2 2 0.927 0.023 0.951 0.916 0.021 0.938
2 3 0.940 0.012 0.962 0.933 0.010 0.951
2 4 0.938 0.023 0.964 0.936 0.018 0.963
2 5 0.903 0.013 0.922 0.909 0.017 0.926
3 1 0.841 0.017 0.863 0.837 0.018 0.864
3 2 0.918 0.019 0.939 0.931 0.010 0.938
3 3 0.865 0.032 0.920 0.889 0.023 0.926
3 4 0.923 0.012 0.937 0.923 0.009 0.938
3 5 0.870 0.014 0.887 0.884 0.022 0.914
4 1 0.799 0.013 0.821 0.790 0.008 0.802
4 2 0.872 0.023 0.895 0.879 0.018 0.901
4 3 0.853 0.022 0.879 0.881 0.020 0.901
4 4 0.891 0.018 0.919 0.909 0.022 0.938
4 5 0.873 0.020 0.910 0.886 0.024 0.926
5 1 0.771 0.039 0.823 0.758 0.032 0.802
5 2 0.838 0.014 0.855 0.857 0.015 0.877
5 3 0.837 0.022 0.861 0.886 0.018 0.914
5 4 0.886 0.014 0.904 0.889 0.016 0.914
5 5 0.860 0.013 0.873 0.896 0.023 0.914
Table 6: The average (avg), standard deviation (std), and best scores for Macro-Accuracy and Accuracy achieved by the Random Forest (RF) classifier over five runs, evaluated for each N-gram size ($N=2$ to $5$) and across multiple data splits, and for fully lowercased versions of the texts.
N-gram Split Macro-Accuracy Accuracy
N no. avg std best avg std best
2 1 0.872 0.017 0.899 0.881 0.018 0.914
2 2 0.879 0.038 0.943 0.921 0.027 0.963
2 3 0.877 0.016 0.899 0.872 0.015 0.889
2 4 0.913 0.028 0.955 0.916 0.026 0.951
2 5 0.903 0.009 0.919 0.886 0.014 0.901
3 1 0.913 0.013 0.926 0.911 0.009 0.926
3 2 0.947 0.016 0.967 0.958 0.006 0.963
3 3 0.944 0.006 0.953 0.936 0.005 0.938
3 4 0.957 0.013 0.974 0.943 0.013 0.963
3 5 0.973 0.013 0.992 0.968 0.013 0.988
4 1 0.929 0.011 0.947 0.926 0.011 0.938
4 2 0.953 0.012 0.974 0.965 0.005 0.975
4 3 0.937 0.009 0.949 0.938 0.008 0.951
4 4 0.970 0.014 0.985 0.958 0.018 0.975
4 5 0.964 0.006 0.972 0.963 0.011 0.975
5 1 0.888 0.041 0.936 0.886 0.038 0.926
5 2 0.950 0.020 0.983 0.968 0.010 0.988
5 3 0.934 0.009 0.944 0.941 0.009 0.951
5 4 0.975 0.010 0.985 0.965 0.009 0.975
5 5 0.979 0.018 1.000 0.980 0.017 1.000
Table 7: The average (avg), standard deviation (std), and best scores for Macro-Accuracy and Accuracy achieved by the Artificial Neural Networks (ANN) classifier over five runs, evaluated for each N-gram size ($N=2$ to $5$) and across multiple data splits, and for the original letter casing of the text.
N-gram Split Macro-Accuracy Accuracy
N no. avg std best avg std best
2 1 0.908 0.019 0.940 0.914 0.021 0.951
2 2 0.868 0.016 0.883 0.906 0.006 0.914
2 3 0.875 0.015 0.901 0.867 0.014 0.889
2 4 0.922 0.017 0.946 0.923 0.016 0.951
2 5 0.895 0.009 0.906 0.886 0.009 0.901
3 1 0.902 0.014 0.924 0.909 0.010 0.926
3 2 0.952 0.013 0.965 0.953 0.009 0.963
3 3 0.940 0.013 0.961 0.938 0.016 0.963
3 4 0.974 0.006 0.984 0.960 0.009 0.975
3 5 0.957 0.015 0.983 0.958 0.010 0.975
4 1 0.911 0.013 0.934 0.914 0.014 0.938
4 2 0.958 0.023 0.974 0.965 0.014 0.975
4 3 0.941 0.016 0.956 0.938 0.014 0.951
4 4 0.977 0.009 0.992 0.965 0.014 0.988
4 5 0.967 0.006 0.972 0.970 0.006 0.975
5 1 0.909 0.020 0.925 0.906 0.017 0.926
5 2 0.967 0.009 0.974 0.968 0.010 0.975
5 3 0.943 0.018 0.960 0.946 0.018 0.963
5 4 0.976 0.007 0.986 0.965 0.005 0.975
5 5 0.974 0.020 1.000 0.975 0.022 1.000
Table 8: The average (avg), standard deviation (std), and best scores for Macro-Accuracy and Accuracy achieved by the Artificial Neural Networks (ANN) classifier over five runs, evaluated for each N-gram size ($N=2$ to $5$) and across multiple data splits, and for fully lowercased versions of the texts.

### v2

@@ Abstract
This study addresses the problem of authorship attribution for Romanian texts using the ROST corpus, a standard benchmark in the field. We systematically evaluate six machine learning techniques—Support Vector Machine (SVM), Logistic Regression (LR), k-Nearest Neighbors (k-NN), Decision Trees (DT), Random Forests (RF), and Artificial Neural Networks (ANN), employing character n-gram features for classification. Among these, the ANN model achieved the highest performance, including perfect classification in four out of fifteen runs when using 5-gram features. These results demonstrate that lightweight, interpretable character n-gram approaches can deliver state-of-the-art accuracy for Romanian authorship attribution, rivaling more complex methods. Our findings highlight the potential of simple stylometric features in resource-constrained or under-studied language settings.
@@ 1 Introduction
  • Preprocessing steps: include text normalization and punctuation handling [2];
@@
  • Computational methods: ranging from traditional statistical techniques (e.g., stylometry, PCA) to machine learning algorithms (e.g., Support Vector Machines, Random Forests), deep learning models (e.g., Recurrent Neural Networks, transformers), and modern large language model (LLM) approaches [6];
@@
  • class imbalance (e.g, in representation of authors or genres),
  • generalizability issues (e.g., on cross-domain or cross-genre texts).
@@
Dataset partitioning into training, testing, and validation subsets also affects performance outcomes, as different splits can yield varying results due to random sampling effects. Thus, researchers commonly employ multiple splits or cross-validation techniques to obtain more reliable performance estimates.
@@ 2.2 N-gram Features in Authorship Attribution
N-grams are contiguous sequences of $N$ items, such as characters, words, or other tokens, extracted from text to capture patterns indicating an author’s unique style. Typically, n-grams are constructed at the character, word, or syntactic level [5].
@@ 2.2.1 Advantages of Character N-grams
  • Robustness: Character n-grams are resilient to infrequent errors such as grammatical mistakes or punctuation slips because their discriminative power derives from frequent, recurring patterns [12].
@@
  • Computational Efficiency: Extracting character n-grams is both straightforward and computationally inexpensive, as it involves direct processing of raw text without the need for complex preprocessing.
@@ 2.3 Computational Methods
A wide range of classification algorithms has been applied to authorship attribution (AA) leveraging character n-gram features [16, 17, 18, 19, 20, 21]. Next, we provide a concise overview of the principal methods utilized in this study:
@@ 2.4.1 Accuracy
Accuracy measures the proportion of correctly classified texts across all classes. While widely used, accuracy can be misleading for imbalanced datasets, as it may overemphasize performance on majority classes.
@@ 3.1 The dataset
  • Avram (2023) leveraged a Romanian-specific BERT model, achieving approximately 87% macro-accuracy [9].
@@ 3.2 Data Preprocessing
  • Punctuation Preservation: All other special characters (punctuation marks) were left unchanged, as patterns in punctuation usage may reflect individual authorial style.
@@ 3.3.2 Vectorization via TF-IDF
  • Cell values = TF-IDF weights for N-gram/document pairs
@@ 3.4.1 Implementation Details
  • Feature Extraction: using TfidfVectorizer from scikit-learn to generate character N-gram feature representations.
@@ 3.4.2 Classification Algorithms
We explored several supervised learning algorithms: Support Vector Machine (SVM), Logistic Regression (LR), k-Nearest Neighbor (k-NN), Decision Trees (DT), Random Forest (RF), and Artificial Neural Networks (ANN).
@@ 3.4.3 Model Evaluation
We use standard metrics including accuracy, macro-accuracy (aka. balanced accuracy or macro-averaged accuracy), and the classification report (precision, recall, F1-score), as provided by scikit-learn.
@@ 3.5 Experimental Procedure
For each of the classification methods described above, we conducted a set of experiments to identify the configuration yielding the best performance. Our investigation included the following key aspects:
  • Parameter Exploration: For k-NN, we evaluated five different values of k to determine the optimal neighborhood size.
  • Robustness to Randomness: To mitigate the effects of stochastic variability inherent in certain models, we repeated the training and evaluation of DT, RF, and ANN five times, each with a different random seed. This procedure allowed us to assess the stability and reliability of the results.
@@ 3.6 Parameter Tuning and Model Selection
k-NN metric: minkowski n_neighbors =$=\{3,5,7,9,11\}$
DT criterion: gini randomstate =$=\{7,17,42,67,101\}$
RF n_estimators: 100 randomstate =$=\{7,17,42,67,101\}$
criterion: gini
ANN activation: relu randomstate =$=\{7,17,42,67,101\}$
@@
Table 1: Summary of machine learning models and their hyperparameters. To ensure that performance differences reflect model characteristics rather than random variation, DT, RF, ANN were each trained and evaluated five times using different random seeds. For k-Nearest Neighbors (k-NN), five different values of the number of neighbors were tested.
@@ 3.7.1 Impact of Letter Casing on Classification Performance
Figure 1 depicts the variation in MAcc differences across all classifiers and n-gram sizes. Overall, lowercase texts tend to yield marginally better results, particularly for $k>3$ and for the k-NN and LR classifiers. However, in most cases, the difference remains below 0.02, which is negligible and likely falls within the range of normal variation due to hyperparameter tuning.
For ANN, the average difference between casing strategies is even smaller, approximately 0.009. Interestingly, at n-gram size $N=5$, the trend reverses slightly, with original-case texts performing better, as further illustrated in Figure 4.
@@ 3.7.2 The Impact of N-gram Size on Classification Performance
A noticeable decline in MAcc is observed for RF (Figure 2e) as the N-gram size increases. Conversely, ANN shows a general upward trend in MAcc with larger N-gram dimensions. For SVM, the highest MAcc is achieved at $k=4$, while k-NN attains its peak performance at $k=2$. The remaining methods do not exhibit a clear monotonic trend in MAcc relative to the N-gram size.
Figure 2: Variation of MAcc across different N-gram sizes ($N=2$ to $5$) for all six classification methods: (A) SVM, (B) LR, (C) k-NN, (D) DT, (E) RF, and (F) ANN. This analysis examines how varying the N-gram size affects the models’ performance. Values shown represent the average of the macro-accuracy scores computed over multiple experimental runs.
@@ 3.7.3 Best Results
ANN achieved the best overall performance, with average accuracy and macro-accuracy (MAcc) values exceeding 0.9, as summarized in Table 2. Tables 3 and 4 present the average accuracy and MAcc for each model across N-gram sizes ($N=2$ to $5$), reported separately for original-case and lowercase texts.
@@
ANN consistently outperformed other models, except at $k=2$, where RF achieved superior results. This trend is illustrated graphically in Figure 3, which visualizes the MAcc values from the aforementioned tables.
For ANN, the highest performance was observed at $k=4$ using original-case letters, and at $k=5$ using lowercase letters—with lowercase slightly outperforming original case by 0.003 in MAcc. Conversely, RF achieved the best results at $k=2$, with average macro-accuracy values of 0.914 (original case) and 0.919 (lowercase). Examining RF results further (Figures 2e ) we observe a slight decrease in average accuracy as $k$ increases, although MAcc values across different $k$ are not clearly separated.
Notably, ANN achieved perfect classification (accuracy = 1.0) in one of the test runs. This raises the question of whether such perfect accuracy reflects a consistent pattern for that particular train-test split, or if it is an isolated occurrence. To address this, we conducted an extended series of 15 experiments using different random seeds:
@@
for both original-case and lowercase texts, for that specific split. As shown in Figure 4, perfect accuracy was reached in four cases: twice for $random\_state=42$ (in both casing conditions) and two additional times for original-case texts. The standard deviation of MAcc across these runs was approximately 0.02 (0.02039 for lowercase, 0.01938 for original case, and 0.02040 overall). No clear pattern emerged relating accuracy variation to the random seed. However, original-case texts tended to perform slightly better on average, with a 0.01 higher MAcc (0.974 vs. 0.962 for lowercase), reversing the trend observed in Tables 3 and 4. Given the small magnitude of these differences, and the reversed trend in the second set of experiments for ANN, we conclude that casing (original vs. lowercase) does not have a significant impact on classification performance.
@@ 3.8 Comparison with Existing Results
Our approach, employing an Artificial Neural Network (ANN), achieved an initial average macro-accuracy of 0.935 across five data splits, with one split yielding perfect classification (macro-accuracy = 1.0). To evaluate the consistency of this result, we extended the analysis to 15 additional runs with randomized seeds (Section 3.7.3). For original-case texts, the expanded trials demonstrated a mean macro-accuracy of 0.974 (with a standard deviation $\sigma\approx 0.02$), aligning with the 0.979 baseline from initial testing on the same split. Perfect classification occurred in four runs: once under lowercase transformation and three times with original casing. Notably, random seed 42 produced perfect classification under both text conditions.
These results demonstrate that our lightweight, character N-gram based approach with ANN matches or surpasses existing benchmarks for Romanian authorship attribution, offering a competitive alternative to more complex models such as BERT or transformer-based approaches.